OpenStudy (anonymous):

what are the solutions for the quadratic equation 5x^2 4x = 12?

OpenStudy (anonymous):

there is no sign between 5x^2 and 4x

OpenStudy (anonymous):

its a minus sign sorry

OpenStudy (anonymous):

5x^2-4X-12=0 =>5x^2-10x+4x-12=0 (x-2)(5x+2)

OpenStudy (anonymous):

sparky, bnb answered your previous question incorrectly. you might want to look into that

OpenStudy (anonymous):

5x^2-4x-12=0 (x-2)(5x+6)=0 x=2 or x=-6/5

OpenStudy (anonymous):

brackett has answered wrong!!!!!!!!!!!

OpenStudy (anonymous):

that is incorrect, nancy

OpenStudy (anonymous):

nancy is doing totally wrong....

OpenStudy (anonymous):

first thing....5x-6=0 does not give u x=6

OpenStudy (anonymous):

(x+2)(5x-6)=5x^2+4x-12.........

OpenStudy (anonymous):

he said it is -4x......

OpenStudy (anonymous):

\[5x ^{2}-4x-12=0\] \[5x ^{2}-10x+6x-12=0\] \[5x(x-2)+6(x-2)=0\] \[(5x+6)(x-2)=0\] x = -6/5 x=2